INFRARED ACTION SPECTROSCOPY OF THE OH-H_{2}O COMPLEX

\underline{BRIDGET A. O’DONNELL}, CRAIG MURRAY, and MARSHA I. LESTER, Department of Chemistry, University of Pennsylvania, Philadelphia, PA 19104-6323.

Hydroxyl radicals (OH) are expected to form strong hydrogen bonds with water (H_{2}O), making complexes of OH with water important for a variety of systems, such as the gaseous environment of the atmosphere, the interface of water and/or ice, and the bulk regions of liquid water, snow, and ice. The study of binary OH-water complexes will shed new light on the intermolecular interaction between the OH radical and water molecule in these environments. In this study, the OH-H_{2}O complex is produced in the gas phase by the association of photolytically generated OH radicals with H_{2}O in the collisional region of a supersonic expansion. Infrared spectroscopy is used to identify a band at 3491 cm^{−1}, displaying a shift of –79 cm^{−1}from its hydroxyl monomer, which is consistent with ab initio predictions of the fundamental OH radical stretch of the complex.^{a} The rotationally structured band profile is modeled using a simulation based on an a-type transition for an open-shell complex with partially quenched orbital angular momentum.^{b} Following infrared excitation of OH-H_{2}O, the OH products of vibrational predissociation are detected by laser-induced fluorescence on the OH A^{2}Σ^{+}–X^{2}Π (1,0) transition. By utilizing information from the OH product state distribution, an upper limit to the binding energy of the complex will be determined.

aY. Xie and H. F. Schaefer, J. Chem. Phys. 98, 8829 (1993).

bM. D. Marshall and M. I. Lester, J. Phys. Chem. B. 109, 8400 (2005).